# Masked Diffusion for Generative Recommendation

### v1

@@ Abstract
Generative recommendation (GR) with semantic IDs (SIDs) has emerged as a promising alternative to traditional recommendation approaches due to its performance gains, capitalization on semantic information provided through language model embeddings, and inference and storage efficiency [rajput2023recommender, deng2025onerec]. Existing GR with SIDs works frame the probability of a sequence of SIDs corresponding to a user’s interaction history using autoregressive modeling. While this has led to impressive next item prediction performances in certain settings, these autoregressive GR with SIDs models suffer from expensive inference due to sequential token-wise decoding, potentially inefficient use of training data and bias towards learning short-context relationships among tokens. Inspired by recent breakthroughs in NLP [sahoo2024simple, shi2025simplified, ou2024absorbing], we propose to instead model and learn the probability of a user’s sequence of SIDs using masked diffusion. Masked diffusion employs discrete masking noise to facilitate learning the sequence distribution, and models the probability of masked tokens as conditionally independent given the unmasked tokens, allowing for parallel decoding of the masked tokens. We demonstrate through thorough experiments that our proposed method consistently outperforms autoregressive modeling. This performance gap is especially pronounced in data-constrained settings and in terms of coarse-grained recall, consistent with our intuitions. Moreover, our approach allows the flexibility of predicting multiple SIDs in parallel during inference while maintaining superior performance to autoregressive modeling. Our code is available at https://github.com/snap-research/MADRec.
@@ 1 Introduction
The GR with SIDs framework has led to such successes in large part by offering a means to incorporate both semantic and collaborative signals in item representations that use an exponentially smaller vocabulary size than traditional sparse IDs [rajput2023recommender, yang2024unifying]. These representations are learned by first executing item SID assignment: assigning tuples of tokens, known as SIDs, to items based on a residual clustering of their text and/or visual feature embeddings extracted from a pretrained language and/or vision model. Then, in the user sequence modeling phase, embeddings of these tokens are trained in concert with a sequential model to learn the probability distribution of the sequences of SIDs corresponding to users’ interaction histories [rajput2023recommender, grid].
Despite tremendous interest in GR with SIDs, the majority of studies focus on design choices around item SID assignment [xiao2025progressive, hua2023index, qu2025tokenrec, wang2024learnable, wang2024eager, liu2024end, tan2024idgenrec, jin2023language, zhu2024cost, kuai2024breaking, houactionpiece], adopting the default autoregressive (AR) modeling from [rajput2023recommender] for the user sequence modeling stage. AR modeling has long been the dominant paradigm in Natural Language Processing (NLP) [rumelhart1985feature, bengio2003neural, sutskever2014sequence], leading to countless breakthroughs in the development of LLMs [vaswani2017attention, radford2018improving, radford2019language, brown2020language]. This makes it a reasonable starting point for modeling SID sequences, and indeed, existing results verify that AR modeling’s efficacy for modeling text sequences translates to modeling users’ SID sequences. However, since AR models model the probability of all future tokens as dependent on all preceding tokens, they must generate tokens sequentially, leading to fundamental upper bounds on inference speed and planning. Moreover, in practice, by training on next token prediction, AR models tend to under-index on global relationships among tokens [khandelwal2018sharp, henighan2020scaling, brunato2023coherent].
@@
In this work, we realize these advantages for GR with SIDs by developing MADRec: MAsked Diffusion over SIDs for Generative Recommendation. To the best of our knowledge, MADRec is the first application of masked diffusion for GR with SIDs, and as such, MADRec prioritizes simplicity and generality in its design. This allows us to make a number of foundational empirical observations regarding the effectiveness of masked diffusion for GR with SIDs:
  • Overall performance. MADRec consistently outperforms standard AR modeling with SIDs (i.e. TIGER [rajput2023recommender]) as well as other GR baselines on several benchmark sequential recommendation datasets. The performance improvement is especially large for coarse-grained recall, suggesting masked diffusion’s strength at learning global item relationships.
  • Data efficiency. Up until a point at which the data becomes too sparse, the performance gap between MADRec and TIGER grows as we shrink the dataset size, supporting the hypothesis that masked diffusion makes better use of limited data.
  • Inference efficiency and flexibility. MADRec can decode multiple tokens in parallel, which allows for flexibly trading off inference performance and efficiency by choosing the number of forward passes (function evaluations) to execute. MADRec already outperforms TIGER with fewer function evaluations, and its performance improves with additional evaluations.
  • Extensibility. MADRec is compatible with auxiliary methods for improving performance in GR. Here we focus on incorporating dense retrieval, motivated by the impressive performance of fusing TIGER with dense retrieval [yang2024unifying, yang2025sparse].
Reproducibility. Our code is available at the following link: https://github.com/snap-research/MADRec.
@@ 2 Generative Recommendation with Semantic IDs
corresponding to the sequence of SIDs of the items interacted with by user $u$.
Existing methods primarily rely on autoregressive modeling, factorizing the probability either by each SID (eq.(1)) or by each item’s full SID sequence (eq.(2)):
@@
3 Proposed Method: MADRec
Recent works have demonstrated the power of masked diffusion models for modeling complex discrete distributions in NLP as well as protein design [lou2024discrete, sahoo2024simple, shi2025simplified, ou2024absorbing]. Motivated by these results and the limitations of AR modeling, we propose MAsked Diffusion over SIDs for Generative Recommendation, i.e. MADRec, a framework that models the probability distribution of SID sequences through masked diffusion.
@@ 3 Proposed Method: MADRec
Figure 1: Overview of training and inference of MADRec. During training, MADRec randomly masks each SID in the SID sequence with a probability $t\sim\textrm{Unif}[0,1]$ and feeds the masked sequence into an encoder-only transformer. The model is then optimized to reconstruct the original values of the masked SIDs using a cross-entropy loss applied at the masked positions (see Eq. (3)). During inference, MADRec begins with all SIDs of the last item replaced by masks. At each inference step, the partially masked sequence is passed through the network to predict values for all masked positions. The model then selectively unmasks a subset of these positions by retaining their predicted values while keeping the remaining positions masked. This iterative process continues until all SIDs are unmasked.
3.1 MADRec Training
@@
After training, MADRec learns the conditional distribution of each masked SID given the remaining sequence. The next section describes the inference procedure and beam search for MADRec.
3.2 MADRec Inference and Beam Search
To describe MADRec inference, consider a user interaction history of item IDs $(i_{1},\ldots,i_{n-1})$ with the corresponding SID sequence $Q=(s_{1}^{1},\ldots,s_{n-1}^{m})$, and assume we want to predict the next item $i_{n}$. To generate a sample using the learned denoising distribution $p_{\theta}(S_{0}^{u}(i)\mid S_{t}^{u})$, MADRec begins by masking the SIDs of the $n$th item, forming $\tilde{S}_{1}=(Q,A_{1})$ where $A_{1}=([\mathrm{M}],\ldots,[\mathrm{M}])$. To transition from $\tilde{S}_{t}=(Q,A_{t})$ to $\tilde{S}_{r}=(Q,A_{r})$ for a lower noise scale $r<t$, the model iteratively unmasks a selected subset of masked tokens, sampling their values from the learned distribution $p_{\theta}(\tilde{S}_{r}(i)\mid\tilde{S}_{t})$.
@@ 3.2 MADRec Inference and Beam Search
Beam Search in MADRec.
Unlike AR models, which decode tokens in a fixed left-to-right order, MADRec can generate SIDs in any order. Therefore, to understand the beam-search, suppose the tokens of the next item are unmasked following a specific order $(k_{1},\ldots,k_{m})$. Beam search in MADRec aims to maximize the probability of the generated item by computing
@@
Beam Search in MADRec with Multi-Token Prediction.
The MADRec framework also supports predicting multiple tokens simultaneously, enabling the generation of the full $m$-length SID sequence with fewer than $m$ sequential function evaluations. Since MADRec models only the conditional probabilities of masked tokens and not the joint probability among the masked tokens, i.e. it assumes the masked tokens are conditionally independent given the unmasked tokens, it approximates the probability of the generated item $p_{\theta}(Q,s_{n}^{k_{1}},\ldots,s_{n}^{k_{m}})$ during beam search.
@@
In other words, we approximate the joint probability of all SIDs $(k^{i}_{1},\ldots,k^{i}_{\alpha_{i}})$ being unmasked at step $i$ by the product of their conditional probabilities. We use this approximation to guide the beam search inference in MADRec while predicting multiple SIDs. Empirically, we find that decoding multiple tokens simultaneously outperforms AR models despite this conditional independence assumption; please see Section 4.4 for more details. In the following section, we demonstrate how to extend MADRec to incorporate an auxiliary method, namely fusing dense and generative retrieval, that was originally developed for AR-based SID modeling.
3.3 Extending MADRec with Dense Retrieval
AR modeling is a dominant paradigm to model the SID sequences in GR. Several prior works have extended it, for instance, by combining SID generation with dense retrieval [yang2024unifying, yang2025sparse], incorporating user preferences [paischer2024preference], or developing methods for long SID generation [hou2025generating]. MADRec presents a novel approach to modeling the probability distribution of SID sequences. We believe that, in principle, MADRec can be integrated with these existing AR extensions. To demonstrate this potential, we integrate MADRec with dense retrieval, inspired by the work of Yang et al. [yang2024unifying].
To unify MADRec’s SID generation capabilities with dense retrieval, we modify the framework to output a dense item embedding in addition to the sequence of SIDs. Specifically, we introduce three key modifications to the MADRec framework:
@@ 3.3 Extending MADRec with Dense Retrieval
  • Masking Strategy: In the original MADRec, each SID is independently masked with a probability proportional to the noise scale $t$. However, since the dense embedding corresponds to the item as a whole, we encourage learning at the item abstraction level by masking all SIDs of an item jointly. Concretely, during training, with a fixed probability $\beta$, we mask all SIDs of an item together, using the same noise scale as the masking probability, and with probability $1-\beta$, we use the masking strategy of the original MADRec. Here, $\beta$ is a hyperparameter.
@@
These modifications extend MADRec to unify SID generation with dense retrieval. We now turn to experiments to examine the performance of MADRec.
@@ 4 Experiments
  • Q1. How does the overall performance of MADRec compare to AR modeling with SIDs and other GR baselines?
  • Q2. How does MADRec’s perform in data-constrained settings?
  • Q3. How does MADRec trade off inference efficiency and performance?
  • Q4. How does MADRec depend on its components?
  • Q5. Does extending MADRec with dense retrieval improve its performance?
@@ 4.1 Experimental Setup
We compare the performance of MADRec against six representative baselines [kang2018self, sun2019bert4rec, yang2023generate, cui2024cadirec, rajput2023recommender, yang2024unifying]:
@@
We begin by presenting our experimental results on the generative recommendation (GR) task, comparing MADRec’s performance with other GR methods across the four aforementioned datasets. The results, shown in Table 1, indicate that our proposed framework performs comparably or better than other GR methods.
@@
Moreover, we further observe in Figure 2 that the performance gap between MADRec and TIGER increases as we increase the retrieval granularity. Here we measure Recall@K for $K\in\{5,10,20,40\}$ on the Beauty and Sports datasets, and notice MADRec’s increasing performance gains with larger $K$, i.e., more coarse-grained retrieval. This suggests that masked diffusion is better able to model global relationships among tokens than AR modeling, as MADRec’s rankings tend to have higher-quality depth than TIGER’s. This is consistent with prior results suggesting that AR modeling over-indexes on local relationships among tokens [khandelwal2018sharp, henighan2020scaling, brunato2023coherent], as well as our intuition that masked diffusion, by training a more diverse set of targets, is better able to model global relationships.
@@
MADRec also outperforms continuous diffusion-based recommendation approaches, namely DreamRec and CaDiRec, as shown in Table 1. Importantly, both DreamRec and CaDiRec require 1000 diffusion inference steps, whereas our method uses only 5 inference steps. This further supports the conclusion of prior works that diffusion with masking noise is more effective for generation in discrete domains than diffusion with Gaussian noise [lou2024discrete, gong2024scaling].
Figure 2: Improved performance gap for coarse-grained retrieval on the Beauty and Sports datasets. The gap in Recall@K between TIGER and MADRec increases as K increases.
Table 1: Performance comparison of MADRec with other GR methods on the Beauty, Sports, Toys, and MovieLens-1M datasets. The best result for each metric is in bold and the second best is underlined.
@@
To better understand MADRec’s effectiveness in data-constrained settings, we evaluate its and TIGER’s performance on increasingly sparsified versions of the Beauty dataset. Specifically, we drop 25%, 37.5%, 50%, 62.5%, and 75% of items from each sequence in the training set, while ensuring each sequence has at least three items. The validation and test sets remain unchanged. We then measure the retained performance by comparing the Recall@5 and NDCG@5 scores of these models with the scores of a model trained without dropping items. The results are shown in Figure 3.
As the percentage of dropped items increases, TIGER’s performance drops faster than MADRec’s. At significantly high drop percentages (62.5% and 75% of items dropped), the performance drop for both methods becomes similar, but this is expected, as dropping close to 100% of the items would result in near-zero performance retention for any method.
Figure 3: Comparison of data efficiency of MADRec and TIGER by dropping 25%, 37.5%, 50%, 67.5% and 75% of items from each sequence in the training set, while maintaining at least three items in each sequence.
@@
Unlike AR methods, MADRec can decode $m$ SIDs of an item with fewer than $m$ sequential function evaluations, albeit with a potential drop in performance. To study the trade-off between the number of function evaluations (NFEs, denoted as $T$ in Section 3.2) and performance, we evaluate MADRec with $2,3,4$ and $5$ NFEs on the Beauty dataset (MADRec uses 5 SIDs per item). We compare the performance using NDCG@5 with AR baselines in Figure 4. Notably, even with only 3 NFEs, MADRec surpasses the performance of TIGER by 13% and LIGER by 4.7%, which both use 4 NFEs.
@@
Intuitively, the benefit of MADRec’s multi-token prediction should become more apparent as the number of items to recommend grows. To evaluate this hypothesis, we adapt the standard leave-one-out evaluation protocol into a leave-two-out protocol: the last two items of each user’s sequence are used for testing, the two preceding items for validation, and the remaining sequence for training. This evaluation protocol also aligns closely with session-wise recommendation tasks [deng2025onerec].
We train and evaluate MADRec and TIGER on the ML-1M dataset using this leave-two-out protocol, choosing ML-1M because of its longer average sequence length. We fix the total number of generated beams to 10. In this setting, each beam represents a pair of items, so the number of candidate recommendations at each position is fewer than 10. After generating 10 session beams, we compute the recall for each predicted item and average the results, which we refer to as the average session recall@10. Figure 4 plots this metric against the number of function evaluations. We observe that MADRec achieves the same performance as TIGER with only 4 NFEs—reducing NFEs by 50% compared to TIGER’s 8.
Figure 4: Next-$k$ item prediction performance vs number of function evaluations (NFEs) during inference for (Left) $k=1$ on Beauty and (Right) $k=2$ on MovieLens-1M. The AR methods (TIGER and LIGER) must decode tokens sequentially, so they always execute $k\;\times$ (# SIDS/item) NFEs. MADRec can decode multiple items in parallel, thereby allows trading off performance and efficiency by tuning the NFEs. Moreover, it already outperforms the AR methods with fewer NFEs.
@@
We next execute several ablation studies to understand the importance of key components of MADRec.
@@
To understand MADRec’s ability to utilize semantic information from the SIDs derived from the item text embeddings, we conduct two complementary experiments on the Beauty dataset in which the provided semantic information is systematically removed. In the first experiment, instead of using SIDs generated from item embeddings, we replace them with randomly assigned tuples of tokens with the same vocabulary as the original SIDs. We refer to this method as MADRec + Random SIDs. In the second experiment, instead of modeling the probability distribution over the SID sequence as in MADRec, we directly model the probability distribution over the item IDs. More specifically, we model the item ID sequence $(i_{1},\ldots,i_{n})$ according to the framework of MADRec in place of the SID sequence $(s_{1}^{1},\ldots,s_{n}^{m})$. We refer to this method as MADRec + Item IDs. We report and compare the performance in Table 2.
We find that replacing the true SIDs computed by residual k-means clustering of semantic item embeddings with random SIDs reduces Recall@10 from 8.15 to 5.53, a drastic reduction. Moreover, using MADRec on item IDs instead of SIDs decreases performance from 8.15 to 6.71. These results demonstrate that MADRec effectively leverages the semantic information contained in the SIDs.
@@
In Table 2, we also compare against BERT4Rec [sun2019bert4rec]. Interestingly, MADRec trained directly on item IDs still outperforms BERT4Rec. BERT4Rec predicts randomly masked item IDs in the interaction history but masks only a fixed fraction of items during training ($t=0.15$), whereas MADRec + Item IDs masks all possible fractions in the interval $[0,1]$, enabling a more effective training regime.
Table 2: Comparison of Recall@K (R@K) and NDCG@K (N@K) for $K\in\{5,10\}$ for different versions of MADRec on the Beauty dataset. Note that BERT4Rec is effectively MADRec + Item IDs with fixed masking ratio.
@@
MADRec w/ Item IDs 4.69 6.71 3.12 3.77
MADRec w/ Random SIDs 3.78 5.53 2.61 3.05
MADRec 5.38 8.15 3.51 4.41

### v2

@@ Abstract
Generative recommendation (GR) with semantic IDs (SIDs) has emerged as a promising alternative to traditional recommendation approaches due to its performance gains, capitalization on semantic information provided through language model embeddings, and inference and storage efficiency [rajput2023recommender, deng2025onerec]. Existing GR with SIDs works frame the probability of a sequence of SIDs corresponding to a user’s interaction history using autoregressive modeling. While this has led to impressive next item prediction performances in certain settings, these autoregressive GR with SIDs models suffer from expensive inference due to sequential token-wise decoding, potentially inefficient use of training data and bias towards learning short-context relationships among tokens. Inspired by recent breakthroughs in NLP [sahoo2024simple, shi2025simplified, ou2024absorbing], we propose to instead model and learn the probability of a user’s sequence of SIDs using masked diffusion. Masked diffusion employs discrete masking noise to facilitate learning the sequence distribution, and models the probability of masked tokens as conditionally independent given the unmasked tokens, allowing for parallel decoding of the masked tokens. We demonstrate through thorough experiments that our proposed method consistently outperforms autoregressive modeling. This performance gap is especially pronounced in data-constrained settings and in terms of coarse-grained recall, consistent with our intuitions. Moreover, our approach allows the flexibility of predicting multiple SIDs in parallel during inference while maintaining superior performance to autoregressive modeling. Our code is available at https://github.com/snap-research/MaskGR.
@@ 1 Introduction
The GR with SIDs framework has led to such successes in large part by offering a means to incorporate both semantic and collaborative signals in item representations that use an exponentially smaller vocabulary size than traditional sparse IDs [rajput2023recommender, yang2024unifying]. These representations are learned by first executing item SID assignment: assigning tuples of tokens, known as SIDs, to items based on a residual clustering of their text and/or visual feature embeddings extracted from a pretrained language and/or vision model. Then, in the user sequence modeling phase, embeddings of these tokens are trained in concert with a sequential model to learn the probability distribution of the sequences of SIDs corresponding to users’ interaction histories [rajput2023recommender, grid]. Despite tremendous interest in GR with SIDs, the majority of studies focus on design choices around item SID assignment [xiao2025progressive, hua2023index, qu2025tokenrec, wang2024learnable, wang2024eager, liu2024end, tan2024idgenrec, jin2023language, zhu2024cost, kuai2024breaking, houactionpiece], adopting the default autoregressive (AR) modeling from [rajput2023recommender] for the user sequence modeling stage.
AR modeling has long been the dominant paradigm in Natural Language Processing (NLP) [rumelhart1985feature, bengio2003neural, sutskever2014sequence], leading to countless breakthroughs in the development of LLMs [vaswani2017attention, radford2018improving, radford2019language, brown2020language]. This makes it a reasonable starting point for modeling SID sequences, and indeed, existing results verify that AR modeling’s efficacy for modeling text sequences translates to modeling users’ SID sequences. However, since AR models model the probability of all future tokens as dependent on all preceding tokens, they must generate tokens sequentially, leading to fundamental upper bounds on inference speed and planning. Moreover, in practice, by training on next token prediction, AR models tend to under-index on global relationships among tokens [khandelwal2018sharp, henighan2020scaling, brunato2023coherent].
@@
In this work, we realize these advantages for GR with SIDs by developing MaskGR: Masked Diffusion over SIDs for Generative Recommendation. To the best of our knowledge, MaskGR is the first application of masked diffusion for GR with SIDs, and as such, MaskGR prioritizes simplicity and generality in its design. This allows us to make a number of foundational empirical observations regarding the effectiveness of masked diffusion for GR with SIDs:
  • Overall performance. MaskGR consistently outperforms standard AR modeling with SIDs (i.e. TIGER [rajput2023recommender]) as well as other GR baselines on several benchmark sequential recommendation datasets. The performance improvement is especially large for coarse-grained recall, suggesting masked diffusion’s strength at learning global item relationships.
  • Data efficiency. Up until a point at which the data becomes too sparse, the performance gap between MaskGR and TIGER grows as we shrink the dataset size, supporting the hypothesis that masked diffusion makes better use of limited data.
  • Inference efficiency and flexibility. MaskGR can decode multiple tokens in parallel, which allows for flexibly trading off inference performance and efficiency by choosing the number of forward passes (function evaluations) to execute. MaskGR already outperforms TIGER with fewer function evaluations, and its performance improves with additional evaluations.
  • Extensibility. MaskGR is compatible with auxiliary methods for improving performance in GR. Here we focus on incorporating dense retrieval, motivated by the impressive performance of fusing TIGER with dense retrieval [yang2024unifying, yang2025sparse].
Reproducibility. Our code is available at the following link: https://github.com/snap-research/MaskGR.
@@ 2 Generative Recommendation with Semantic IDs
corresponding to the sequence of SIDs of the items interacted with by user $u$. Existing methods primarily rely on autoregressive modeling, factorizing the probability either by each SID (eq.(1)) or by each item’s full SID sequence (eq.(2)):
@@
3 Proposed Method: MaskGR
Recent works have demonstrated the power of masked diffusion models for modeling complex discrete distributions in NLP as well as protein design [lou2024discrete, sahoo2024simple, shi2025simplified, ou2024absorbing]. Motivated by these results and the limitations of AR modeling, we propose Masked Diffusion over SIDs for Generative Recommendation, i.e. MaskGR, a framework that models the probability distribution of SID sequences through masked diffusion.
@@ 3 Proposed Method: MaskGR
Figure 1: Overview of training and inference of MaskGR. During training, MaskGR randomly masks each SID in the SID sequence with a probability $t\sim\textrm{Unif}[0,1]$ and feeds the masked sequence into an encoder-only transformer. The model is then optimized to reconstruct the original values of the masked SIDs using a cross-entropy loss applied at the masked positions (see Eq. (3)). During inference, MaskGR begins with all SIDs of the last item replaced by masks. At each inference step, the partially masked sequence is passed through the network to predict values for all masked positions. The model then selectively unmasks a subset of these positions by retaining their predicted values while keeping the remaining positions masked. This iterative process continues until all SIDs are unmasked.
3.1 MaskGR Training
@@
After training, MaskGR learns the conditional distribution of each masked SID given the remaining sequence. The next section describes the inference procedure and beam search for MaskGR.
3.2 MaskGR Inference and Beam Search
To describe MaskGR inference, consider a user interaction history of item IDs $(i_{1},\ldots,i_{n-1})$ with the corresponding SID sequence $Q=(s_{1}^{1},\ldots,s_{n-1}^{m})$, and assume we want to predict the next item $i_{n}$. To generate a sample using the learned denoising distribution $p_{\theta}(S_{0}^{u}(i)\mid S_{t}^{u})$, MaskGR begins by masking the SIDs of the $n$th item, forming $\tilde{S}_{1}=(Q,A_{1})$ where $A_{1}=([\mathrm{M}],\ldots,[\mathrm{M}])$. To transition from $\tilde{S}_{t}=(Q,A_{t})$ to $\tilde{S}_{r}=(Q,A_{r})$ for a lower noise scale $r<t$, the model iteratively unmasks a selected subset of masked tokens, sampling their values from the learned distribution $p_{\theta}(\tilde{S}_{r}(i)\mid\tilde{S}_{t})$.
@@ 3.2 MaskGR Inference and Beam Search
Beam Search in MaskGR.
Unlike AR models, which decode tokens in a fixed left-to-right order, MaskGR can generate SIDs in any order. Therefore, to understand the beam-search, suppose the tokens of the next item are unmasked following a specific order $(k_{1},\ldots,k_{m})$. Beam search in MaskGR aims to maximize the probability of the generated item by computing
@@
Beam Search in MaskGR with Multi-Token Prediction.
The MaskGR framework also supports predicting multiple tokens simultaneously, enabling the generation of the full $m$-length SID sequence with fewer than $m$ sequential function evaluations. Since MaskGR models only the conditional probabilities of masked tokens and not the joint probability among the masked tokens, i.e. it assumes the masked tokens are conditionally independent given the unmasked tokens, it approximates the probability of the generated item $p_{\theta}(Q,s_{n}^{k_{1}},\ldots,s_{n}^{k_{m}})$ during beam search.
@@
In other words, we approximate the joint probability of all SIDs $(k^{i}_{1},\ldots,k^{i}_{\alpha_{i}})$ being unmasked at step $i$ by the product of their conditional probabilities. We use this approximation to guide the beam search inference in MaskGR while predicting multiple SIDs. Empirically, we find that decoding multiple tokens simultaneously outperforms AR models despite this conditional independence assumption; please see Section 4.4 for more details. In the following section, we demonstrate how to extend MaskGR to incorporate an auxiliary method, namely fusing dense and generative retrieval, that was originally developed for AR-based SID modeling.
3.3 Extending MaskGR with Dense Retrieval
AR modeling is a dominant paradigm to model the SID sequences in GR. Several prior works have extended it, for instance, by combining SID generation with dense retrieval [yang2024unifying, yang2025sparse], incorporating user preferences [paischer2024preference], or developing methods for long SID generation [hou2025generating]. MaskGR presents a novel approach to modeling the probability distribution of SID sequences. We believe that, in principle, MaskGR can be integrated with these existing AR extensions. To demonstrate this potential, we integrate MaskGR with dense retrieval, inspired by the work of Yang et al. [yang2024unifying].
To unify MaskGR’s SID generation capabilities with dense retrieval, we modify the framework to output a dense item embedding in addition to the sequence of SIDs. Specifically, we introduce three key modifications to the MaskGR framework:
@@ 3.3 Extending MaskGR with Dense Retrieval
  • Masking Strategy: In the original MaskGR, each SID is independently masked with a probability proportional to the noise scale $t$. However, since the dense embedding corresponds to the item as a whole, we encourage learning at the item abstraction level by masking all SIDs of an item jointly. Concretely, during training, with a fixed probability $\beta$, we mask all SIDs of an item together, using the same noise scale as the masking probability, and with probability $1-\beta$, we use the masking strategy of the original MaskGR. Here, $\beta$ is a hyperparameter.
@@
These modifications extend MaskGR to unify SID generation with dense retrieval. We now turn to experiments to examine the performance of MaskGR.
@@ 4 Experiments
  • Q1. How does the overall performance of MaskGR compare to AR modeling with SIDs and other GR baselines?
  • Q2. How does MaskGR’s perform in data-constrained settings?
  • Q3. How does MaskGR trade off inference efficiency and performance?
  • Q4. How does MaskGR depend on its components?
  • Q5. Does extending MaskGR with dense retrieval improve its performance?
@@ 4.1 Experimental Setup
We compare the performance of MaskGR against six representative baselines [kang2018self, sun2019bert4rec, yang2023generate, cui2024cadirec, rajput2023recommender, yang2024unifying]:
@@
We begin by presenting our experimental results on the generative recommendation (GR) task, comparing MaskGR’s performance with other GR methods across the four aforementioned datasets. The results, shown in Table 1, indicate that our proposed framework performs comparably or better than other GR methods.
@@
Moreover, we further observe in Figure 2 that the performance gap between MaskGR and TIGER increases as we increase the retrieval granularity. Here we measure Recall@K for $K\in\{5,10,20,40\}$ on the Beauty and Sports datasets, and notice MaskGR’s increasing performance gains with larger $K$, i.e., more coarse-grained retrieval. This suggests that masked diffusion is better able to model global relationships among tokens than AR modeling, as MaskGR’s rankings tend to have higher-quality depth than TIGER’s. This is consistent with prior results suggesting that AR modeling over-indexes on local relationships among tokens [khandelwal2018sharp, henighan2020scaling, brunato2023coherent], as well as our intuition that masked diffusion, by training a more diverse set of targets, is better able to model global relationships.
@@
MaskGR also outperforms continuous diffusion-based recommendation approaches, namely DreamRec and CaDiRec, as shown in Table 1. Importantly, both DreamRec and CaDiRec require 1000 diffusion inference steps, whereas our method uses only 5 inference steps. This further supports the conclusion of prior works that diffusion with masking noise is more effective for generation in discrete domains than diffusion with Gaussian noise [lou2024discrete, gong2024scaling].
Figure 2: Improved performance gap for coarse-grained retrieval on the Beauty and Sports datasets. The gap in Recall@K between TIGER and MaskGR increases as K increases.
Table 1: Performance comparison of MaskGR with other GR methods on the Beauty, Sports, Toys, and MovieLens-1M datasets. The best result for each metric is in bold and the second best is underlined.
@@
To better understand MaskGR’s effectiveness in data-constrained settings, we evaluate its and TIGER’s performance on increasingly sparsified versions of the Beauty dataset. Specifically, we drop 25%, 37.5%, 50%, 62.5%, and 75% of items from each sequence in the training set, while ensuring each sequence has at least three items. The validation and test sets remain unchanged. We then measure the retained performance by comparing the Recall@5 and NDCG@5 scores of these models with the scores of a model trained without dropping items. The results are shown in Figure 3.
As the percentage of dropped items increases, TIGER’s performance drops faster than MaskGR’s. At significantly high drop percentages (62.5% and 75% of items dropped), the performance drop for both methods becomes similar, but this is expected, as dropping close to 100% of the items would result in near-zero performance retention for any method.
Figure 3: Comparison of data efficiency of MaskGR and TIGER by dropping 25%, 37.5%, 50%, 67.5% and 75% of items from each sequence in the training set, while maintaining at least three items in each sequence.
@@
Unlike AR methods, MaskGR can decode $m$ SIDs of an item with fewer than $m$ sequential function evaluations, albeit with a potential drop in performance. To study the trade-off between the number of function evaluations (NFEs, denoted as $T$ in Section 3.2) and performance, we evaluate MaskGR with $2,3,4$ and $5$ NFEs on the Beauty dataset (MaskGR uses 5 SIDs per item). We compare the performance using NDCG@5 with AR baselines in Figure 4. Notably, even with only 3 NFEs, MaskGR surpasses the performance of TIGER by 13% and LIGER by 4.7%, which both use 4 NFEs.
@@
Intuitively, the benefit of MaskGR’s multi-token prediction should become more apparent as the number of items to recommend grows. To evaluate this hypothesis, we adapt the standard leave-one-out evaluation protocol into a leave-two-out protocol: the last two items of each user’s sequence are used for testing, the two preceding items for validation, and the remaining sequence for training. This evaluation protocol also aligns closely with session-wise recommendation tasks [deng2025onerec].
We train and evaluate MaskGR and TIGER on the ML-1M dataset using this leave-two-out protocol, choosing ML-1M because of its longer average sequence length. We fix the total number of generated beams to 10. In this setting, each beam represents a pair of items, so the number of candidate recommendations at each position is fewer than 10. After generating 10 session beams, we compute the recall for each predicted item and average the results, which we refer to as the average session recall@10. Figure 4 plots this metric against the number of function evaluations. We observe that MaskGR achieves the same performance as TIGER with only 4 NFEs—reducing NFEs by 50% compared to TIGER’s 8.
Figure 4: Next-$k$ item prediction performance vs number of function evaluations (NFEs) during inference for (Left) $k=1$ on Beauty and (Right) $k=2$ on MovieLens-1M. The AR methods (TIGER and LIGER) must decode tokens sequentially, so they always execute $k\;\times$ (# SIDS/item) NFEs. MaskGR can decode multiple items in parallel, thereby allows trading off performance and efficiency by tuning the NFEs. Moreover, it already outperforms the AR methods with fewer NFEs.
@@
We next execute several ablation studies to understand the importance of key components of MaskGR.
@@
To understand MaskGR’s ability to utilize semantic information from the SIDs derived from the item text embeddings, we conduct two complementary experiments on the Beauty dataset in which the provided semantic information is systematically removed. In the first experiment, instead of using SIDs generated from item embeddings, we replace them with randomly assigned tuples of tokens with the same vocabulary as the original SIDs. We refer to this method as MaskGR + Random SIDs. In the second experiment, instead of modeling the probability distribution over the SID sequence as in MaskGR, we directly model the probability distribution over the item IDs. More specifically, we model the item ID sequence $(i_{1},\ldots,i_{n})$ according to the framework of MaskGR in place of the SID sequence $(s_{1}^{1},\ldots,s_{n}^{m})$. We refer to this method as MaskGR + Item IDs. We report and compare the performance in Table 2.
We find that replacing the true SIDs computed by residual k-means clustering of semantic item embeddings with random SIDs reduces Recall@10 from 8.15 to 5.53, a drastic reduction. Moreover, using MaskGR on item IDs instead of SIDs decreases performance from 8.15 to 6.71. These results demonstrate that MaskGR effectively leverages the semantic information contained in the SIDs.
@@
In Table 2, we also compare against BERT4Rec [sun2019bert4rec]. Interestingly, MaskGR trained directly on item IDs still outperforms BERT4Rec. BERT4Rec predicts randomly masked item IDs in the interaction history but masks only a fixed fraction of items during training ($t=0.15$), whereas MaskGR + Item IDs masks all possible fractions in the interval $[0,1]$, enabling a more effective training regime.
Table 2: Comparison of Recall@K (R@K) and NDCG@K (N@K) for $K\in\{5,10\}$ for different versions of MaskGR on the Beauty dataset. Note that BERT4Rec is effectively MaskGR + Item IDs with fixed masking ratio.
@@
MaskGR w/ Item IDs 4.69 6.71 3.12 3.77
MaskGR w/ Random SIDs 3.78 5.53 2.61 3.05
MaskGR 5.38 8.15 3.51 4.41